hartnn (hartnn):

2,4,8,16,24,36,64,...,... Find the next 2 numbers in the series and write a function to model it.

hartnn (hartnn):

i tried, its not a polynomial function...

OpenStudy (anonymous):

its a geometric series...

OpenStudy (anonymous):

Seems geometric to me, which would follow the general structure of \[\Large a_n=a_1q^{n-1} \]

hartnn (hartnn):

nopes, it isn't...

OpenStudy (anonymous):

true, the 36

OpenStudy (anonymous):

oh your right.. lol fooled two of us XD

OpenStudy (anonymous):

well it shouldn't be too far away, seems just like it changes depending on the \[ q \]

hartnn (hartnn):

as long as you can model it into a function, yes.

OpenStudy (anonymous):

hatnn

OpenStudy (anonymous):

are u viewing my question

OpenStudy (anonymous):

hartnn

hartnn (hartnn):

i also tried \(2^n \pm something\) but no use...

OpenStudy (sirm3d):

\[\Large a_n=\begin{cases}2^{\frac{n+1}{2}}\left(\frac{n+1}{2}\right)&,n \text{ odd}\\n^2&,n \text{ even}\end{cases}\]

hartnn (hartnn):

nice! but can't that be written in single line ?

OpenStudy (sirm3d):

the odd-terms is semi exponential, the even-terms is quadratic. i guess there's no way to write it as a single expression

OpenStudy (anonymous):

what method did you use sirm3d?

OpenStudy (anonymous):

does it start at n=0,1...?

OpenStudy (sirm3d):

odd-terms, even-terms. solve each sequence separately.

hartnn (hartnn):

i tried to do it separately for odd and even terms...but could not figure it out for odd terms...so thanks!

OpenStudy (sirm3d):

yw

hartnn (hartnn):

any paticular method to follow ?

OpenStudy (sirm3d):

i factor the 2's completely in the \odd-terms. the other factors form an arithmetic sequence.

hartnn (hartnn):

good thought :)

OpenStudy (anonymous):

2^(3/2)*3/2=4.2426 for n=3????

OpenStudy (sirm3d):

it's \((n+1)\) in the numerator, not \(n\)

OpenStudy (anonymous):

OIC, great work :)